4mg of NACL was weighed and topped up to 50ml.

Calculate the normality of Cl-




1
Expert's answer
2021-06-28T08:11:22-0400

We have to find no of moles first

So

n=0.004/58.5

n=6.83×10^-5

Normality =n/v

=6.83×10^-5/0.05

Normality =1.37×10^-3